p(x)=x^3-x+1 c=1/2 solve by using synthetic division and the remainder theorem
Answer by stanbon(75887)

p(x)=x^3-x+1 c=1/2 solve by using synthetic division and the remainder theorem
----------
1/2)....1....0....-1....1
........1....1/2..-3/4..|..1/8
------
quotient:: x^2 + (1/2)x - (3/4)
remainder: 1/8
------------------
Cheers,
Stan H.
--------------